Two small objects each of mass m = 0.2 kg are conn
Answer #1

Mass of the object, m- 0.2 kg Length of the rod, d-1.5 m Magnitude of the velocity, หุ่-36 m/s, v-50 m/s Magnitude of the extTotal angular momentum about the point A, = myri sin g +myr, sin a Consider the diagram, rsing=dth =1.5+0.4 = 1.9 m 0.4 m SubTranslational angular momentum, The distance is given by, 1.5 0.4 Substituting, 1.15x0.4x43 19.78kgm/s (along -ve z axis)Now equation (1) implies -17.68-L +-19.78 Hence the rotational angular momentum, L,,-1978-1768 =2.1kgm /s